OpenStudy (anonymous):

write the slope-intercept equation for the line containing the given pair of points:(-3,4) and (3,1)

OpenStudy (amistre64):

-3 4 -3-1 ----- -6 3; m = -3/6 = -1/2 y = -1/2 x +1(-3)/2 +4 y = -1/2 x -3/2 + 8/2 y = -1/2 x +5/2

OpenStudy (anonymous):

Thank you amistre64 for your response. Robbie